# Interpretable User Satisfaction Estimation for Conversational Systems with Large Language Models

### v1

@@ Abstract
Accurate and interpretable user satisfaction estimation (USE) is critical for understanding, evaluating, and continuously improving conversational systems. Users express their satisfaction or dissatisfaction with diverse conversational patterns in both general-purpose (ChatGPT and Bing Copilot) and task-oriented (customer service chatbot) conversational systems. Existing approaches based on featurized ML models or text embeddings fall short in extracting generalizable patterns and are hard to interpret. In this work, we show that LLMs can extract interpretable signals of user satisfaction from their natural language utterances more effectively than embedding-based approaches. Moreover, an LLM can be tailored for USE via an iterative prompting framework using supervision from labeled examples. The resulting method, Supervised Prompting for User satisfaction Rubrics (SPUR), not only has higher accuracy but is more interpretable as it scores user satisfaction via learned rubrics with a detailed breakdown.
@@
Ying-Chun ${}^{*{\ddagger}}$, Jennifer ${}^{*{\dagger}}$, Jack W. ${}^{*{\dagger}}$, Longqi ${}^{*{\dagger}}$,thanks: $*{}^{*}$hese corresponding authors contributed equally to this work. Email: lin915@purdue.edu, jenneville@microsoft.com, jstokes@microsoft.com, longqi.yang@microsoft.com Tara ${}^{{\dagger}}$, Mengting ${}^{{\dagger}}$, Scott ${}^{{\dagger}}$, Siddharth ${}^{{\dagger}}$, Reid ${}^{{\dagger}}$, Xiaofeng ${}^{{\dagger}}$, Deepak ${}^{{\dagger}}$, Sujay Kumar ${}^{{\dagger}}$, Xia ${}^{{\dagger}}$, Georg ${}^{{\dagger}}$, Saurabh ${}^{{\dagger}}$, Brent ${}^{{\dagger}}$, Jaime ${}^{{\dagger}}$$^{{\dagger}}$icrosoft Corporation, ${}^{{\ddagger}}$urdue University
@@ 1 Introduction
General-purpose conversational systems such as ChatGPT and Bing Copilot are revolutionizing how people live and work. Understanding when and why users are satisfied or dissatisfied is critical for the continuous improvement of these systems. It helps system developers identify areas of improvements, conduct effective A/B experiments, and optimize underlying models. Unsurprisingly, developing machine learning models for User Satisfaction Estimation (USE) (Hu et al., 2023; Kachuee et al., 2021a; Song et al., 2019; Bodigutla et al., 2019, 2020) has captured significant attention from the research community.
@@
When estimating user satisfaction, simply classifying that a user is satisfied or dissatisfied is insufficient. Understanding the reason why a user is satisfied or dissatisfied is just as valuable (Figure 1), e.g., frequent query reformulation presents opportunities for prompt recommendation, and conversations where users explicitly correct a bot’s mistakes can suggest examples for model alignment. However, most of the existing work has focused on improving the classification accuracy but overlooked the interpretability. For example, representation learning-based approaches (Song et al., 2023; Deng et al., 2022; Ye Fanghua, 2023) are relatively opaque due to their use of neural models (e.g., embeddings) and thus offer little insight into conversational patterns that indicate satisfaction/dissatisfaction. Similar limitations apply to reward models for training LLMs, e.g., RLHF (Christiano et al., 2017) and RLAIF (Bai et al., 2022). In this case, the learned model produces a continuous “reward” score that aims to distinguish outputs that a human prefers without explaining why a conversation has a higher score than others. To our knowledge, these reward models have not been directly used for USE, but we treat it as a baseline due to their ability to rank outputs with respect to human preferences.
Some prior work addressed the interpretation needs of USE by a featurized ML model. Examples include Walker et al. (1997), which evaluated user satisfaction based on human-annotated features in terms of task success and dialogue costs, and Bodigutla et al. (2019) that proposed domain-independent features that evaluate response quality. However, the growth of LLM-based conversational systems (e.g., ChatGPT, Bing Copilot) means user queries in conversational systems may now reflect manually crafted domains and intents (e.g., task-oriented, QA, chitchat). As such, approaches based on domain-specific features have limited generalizability to these diverse conversational patterns (Deriu et al., 2021).
In this work, we make the key observation that LLMs can achieve both high classification accuracy and fine-grained interpretability at the same time – through their ability to reason about user conversational patterns and identify salient patterns that generalize and produce accurate predictions. We propose Supervised Prompting for User satisfaction Rubrics (SPUR). We consider a few-shot scenario, where a small number of training examples are available, and develop a supervised, iterative prompting framework that uses an LLM to (1) extract signals of satisfaction from user utterances in a labeled training set, (2) summarize the reasons into rubrics for identifying satisfaction/dissatisfaction conversational patterns, and (3) apply the rubrics to predict satisfaction labels on unseen conversations.
@@
  • We use knowledge distillation to scale the application of the learned rubrics with knowledge distillation and show that the rubrics can continuously improve performance on USE as more training data is available.
@@ 2 Problem Definition and Related Work
Problem Definition. Let a conversation $C$ from session $i$ and consisting of $t$ interaction turns of user-agent utterances be $C_{i}=[U_{1},A_{1},\ldots,U_{t},A_{t}]$. Here $U_{t}$ refers to a user utterance and $A_{t}$ refers to an AI agent utterance. The user-agent utterances $C_{i}$ typically consist of multiple turns, e.g. $t>1$. The conversation also has an overall user satisfaction label $y_{i}\in[-1,+1]$.
Our goal is to learn a function $f:C\rightarrow y$ to accurately predict the satisfaction label of unseen conversations and explain why the label is predicted. In multi-turn conversational sessions, a user can convey their satisfaction (or dissatisfaction) explicitly in their utterances or implicitly through their behavioral interactions with the agent. We refer to these satisfaction/dissatisfaction conversational patterns as SAT/DSAT patterns. Let $\mathcal{S}=\{s_{1},s_{2},\cdots,s_{\infty}\}$ and $\mathcal{D}=\{d_{1},d_{2},\cdots,d_{\infty}\}$ be the set of all interpretable SAT and DSAT patterns respectively. We assume these are latent and unknown. The goal is to identify a subset of SAT/DSAT patterns ($\mathcal{S}_{s}\subset\mathcal{S},\mathcal{D}_{s}\subset\mathcal{D}$) that summarize the conversation enough to accurately predict its label: $P(y|C)\approx P\Big{(}y\>\Big{|}\>\mathcal{S}_{s}(C),\mathcal{D}_{s}(C)\Big{)}$.
@@
Related Work. Numerous prior research studies have examined User Satisfaction Evaluation (USE) through the lenses of sentiment analysis (Song et al., 2023, 2019), content analysis (Walker et al., 1997; Sun et al., 2021), and response quality assessment (Schmitt and Ultes, 2015; Bodigutla et al., 2019). While analyzing user sentiment distribution in a dialogue session can enhance the model’s USE capabilities, it is important to note that sentiment analysis is not equivalent to USE (Song et al., 2023). Another common approach involves content analysis, which typically necessitates the employment of human annotators to evaluate interaction quality in a dialogue session (Schmitt and Ultes, 2015; Bodigutla et al., 2019). Subsequently, a classifier is trained to predict user satisfaction based on the features extracted from the annotation process.
@@
Recently, Large Language Models (LLMs) revolutionized the traditional learning framework (Kojima et al., 2022; Wei et al., 2022), especially in the natural language processing (NLP) area. LLMs have achieved performance comparable to supervised baselines or state-of-the-art results across various NLP tasks with In-Context Learning (ICL). By providing a few examples or hints (Lampinen et al., 2022; Sun et al., 2023) and simple reasoning process (Kojima et al., 2022; Wei et al., 2022), LLMs can have significant performance boosts in NLP tasks. Hu et al. (2023) further use LLMs as a user simulator for USE and adopt the user simulator into RLAIF (Bai et al., 2022) for fine-tuning the existing LLM models. For USE with zero-shot prompting (Kojima et al., 2022; Hu et al., 2023), instructions provided by a human may not fit the actual conversation patterns in the data and hence introduce bias. For few-shot prompting Lampinen et al. (2022); Sun et al. (2023), the provided examples are not enough to describe the full distribution of the conversational patterns, and this results in inaccuracies for USE.
@@ 3 SPUR
We propose SPUR for interpretable User Satisfaction Estimation (USE) given conversations $C$ from a conversational system. Our approach follows the three-phase prompting strategy depicted in Figure 2: Supervised Extraction, Rubric Summarization, and User Satisfaction Estimation. Due to the multi-turn and general-purpose nature in such conversational systems, users demonstrate a variety of response patterns when expressing satisfaction or dissatisfaction. Our three-phase approach of SPUR is essential for enhancing its generalization and accuracy for User Satisfaction Estimation (USE). Through Supervised Extraction, SPUR improves generalization by capturing the diverse conversational patterns for USE. In the Rubric Summarization stage, the LLM explicitly identifies prominent SAT and DSAT patterns in the training set $\mathcal{C}_{train}=\{C_{1},C_{2},\cdots,C_{N}\}$, which are annotated with click feedback. Finally, SPUR uses the learned rubrics generated from the previous stage to score user satisfaction on unlabeled conversations. For the ease of understanding, we use mathematical definitions to approximate the process of SPUR in the following three sections.
@@ 3.1 Supervised Extraction
The first step of our framework is Supervised Extraction—where we use a prompt to obtain meaningful and interpretable SAT/DSAT patterns from GPT-4, which has an exceptional ability for natural language understanding and reasoning (Ye and Durrett, 2023; Huang et al., 2023; Kojima et al., 2022). Given a conversation $C_{i}$ with its user satisfaction label $y_{i}=+1$, how the user expresses satisfaction in $C_{i}$ can be formulated as: $\widehat{\mathbf{s}}_{i}\approx\underset{s\in\mathcal{S}}{\operatorname{arg\,max_{k}}}\;P(\mathcal{S}|C_{i},y_{i}=+1)$, where $\mathcal{S}=\{s_{1},s_{2},\cdots,s_{\infty}\}$ is the set of all possible SAT patterns.
The goal is to identify the top-k potential patterns $\widehat{\mathbf{s}}_{i}=\{s_{1},s_{2},\cdots,s_{k}\}\subset\mathcal{S}$ that are exhibited in $C_{i}$ relevant to satisfaction expression. Similarly, $\widehat{\mathbf{d}}_{i}\approx\underset{d\in\mathcal{D}}{\operatorname{arg\,max_{k}}}\;P(\mathcal{D}|C_{i},y_{i}=-1)$, where $\mathcal{D}=\{d_{1},d_{2},\cdots,d_{\infty}\}$ is the set of all possible DSAT patterns.
The prompt for generating the possible $\widehat{\mathbf{s}}$ or $\widehat{\mathbf{d}}$ patterns from $\mathcal{C}_{train}$ is provided in Appendix A.1. In our prompt, we specifically require GPT-4 to restrict $k\leq 3$ for each $C_{i}$. The prompt for DSAT patterns is similar; we only replace “satisfaction” with “dissatisfaction” in the instructions. For the ease of discussion in the next section, let $\widehat{\mathcal{S}}=\{\widehat{s}_{1},\cdots,\widehat{s}_{N}\}$ denote all the SAT patterns derived from Supervised Extraction and $\widehat{\mathcal{D}}=\{\widehat{d}_{1},\cdots,\widehat{d}_{N}\}$ are all the DSAT patterns.
@@ 3.2 Rubric Summarization
There are two benefits to utilizing the LLM generated satisfaction and dissatisfaction rubrics from this iterative process. First, the rubrics are developed in a supervised manner from the set of training conversations, $\mathcal{C}_{train}$, thereby ensuring that prominent (and thus predictive) SAT and DSAT patterns in the distribution are identified. As a result, the generated rubrics provide a clear guideline for GPT-4 to estimate user satisfaction accurately. Second, the rubrics are generated from more examples than can fit in a single context window. As such, Rubric Summarization improves the generalization for GPT-4 in terms of in-context learning.
@@ 4 Evaluation
Baselines. We compare SPUR with two LLM-based methods, including ZeroShot and FewShot, and three embedding-based methods, Linear Regression, USDA (Deng et al., 2022) and ASAP (Ye Fanghua, 2023). The detailed descriptions of these models are provided in Appendix B.
Dataset. We use four datasets to evaluate the performance of the compared methods. Bing Copilot is a general-purpose and multilingual conversational system, and this dataset includes 50K fully de-identified conversations. [^1] All personal, private or sensitive information were scrubbed and masked before the conversations were used for this research. The access to the dataset is strictly limited to the authors who conducted hands-on analysis and model development. MWOZ, SGD and ReDial are three task-oriented, English conversational systems, and they have 1,155, 1,638 and 1,387 conversations, respectively. We employ the user satisfaction labels provided by Sun et al. (2021) for these task-oriented datasets. Because MWOZ (Eric et al., 2020), SGD (Rastogi et al., 2020), and ReDial (Siro et al., 2022) are labeled turn-by-turn, we further process these labels into a label to represent overall satisfaction of the whole conversation. The preprocessing details are described in Appendix C.
Ethics. As part of the production process, the Bing Copilot data is de-identified, and each conversation is formed by aggregating turns based on a unique conversation ID. Thus, none of the researchers who analyzed the data are able to recover and identify the conversations from any individual user. In addition, this research study was reviewed and approved by representatives from our institutional review board (IRB), as well as our ethics and security teams. No formal IRB certificate was required as we did not conduct human studies for this paper.
Table 1: Accuracy (Acc), Precision (P), Recall (R), and F1 Score (F1) on USE with small training set sizes. The training sizes are shown besides the name of each dataset. The testing size is $80\%$ of the data. The best scores are in bold face.

### v2

@@ Abstract
Accurate and interpretable user satisfaction estimation (USE) is critical for understanding, evaluating, and continuously improving conversational systems. Users express their satisfaction or dissatisfaction with diverse conversational patterns in both general-purpose (ChatGPT and Bing Copilot) and task-oriented (customer service chatbot) conversational systems. Existing approaches based on featurized ML models or text embeddings fall short in extracting generalizable patterns and are hard to interpret. In this work, we show that LLMs can extract interpretable signals of user satisfaction from their natural language utterances more effectively than embedding-based approaches. Moreover, an LLM can be tailored for USE via an iterative prompting framework using supervision from labeled examples. Our proposed method, Supervised Prompting for User satisfaction Rubrics (SPUR), not only has higher accuracy but is more interpretable as it scores user satisfaction via learned rubrics with a detailed breakdown.
@@
Ying-Chun Lin∗‡, Jennifer Neville∗†, Jack W. Stokes∗†, Longqi Yang∗†,thanks: These corresponding authors contributed equally to this work. Email: lin915@purdue.edu, jenneville@microsoft.com, jstokes@microsoft.com, longqi.yang@microsoft.com Tara Safavi, Mengting Wan, Scott Counts, Siddharth Suri, Reid Andersen, Xiaofeng Xu, Deepak Gupta, Sujay Kumar Jauhar, Xia Song, Georg Buscher, Saurabh Tiwary, Brent Hecht, Jaime Teevan Microsoft Corporation, Purdue University
@@ 1 Introduction
General-purpose conversational systems such as ChatGPT and Copilot are revolutionizing how people live and work. Understanding when and why users are satisfied or dissatisfied is critical for the continuous improvement of these systems. It helps system developers identify areas of improvements, conduct effective A/B experiments, and optimize underlying models. Unsurprisingly, developing machine learning models for User Satisfaction Estimation (USE) (Hu et al., 2023; Kachuee et al., 2021a; Song et al., 2019; Bodigutla et al., 2019, 2020) has captured significant attention from the research community.
@@
When estimating user satisfaction, simply classifying that a user is satisfied or dissatisfied is insufficient. Understanding the reason why a user is satisfied or dissatisfied is just as valuable. For example, frequent query reformulation presents opportunities for prompt recommendation and conversations where users explicitly correct a bot’s mistakes can suggest examples for model alignment. See Figure 1 for an illustration. However, most existing work has focused on improving classification accuracy and has overlooked interpretability. Representation learning-based approaches (Song et al., 2023; Deng et al., 2022; Ye Fanghua, 2023) are relatively opaque due to their use of neural models (e.g., embeddings) and thus offer little insight into conversational patterns that indicate satisfaction/dissatisfaction. Similar limitations apply to reward models for training LLMs, e.g., RLHF (Christiano et al., 2017) and RLAIF (Bai et al., 2022). In this case, the learned model produces a continuous “reward” score that aims to distinguish outputs that a human prefers without explaining why a conversation has a higher score than others. To our knowledge, these reward models have not been directly used for USE, but we treat it as a baseline due to their ability to rank outputs with respect to human preferences.
Some prior work addressed the interpretation needs of USE via featurized ML models. Examples include Walker et al. (1997), which evaluated user satisfaction based on human-annotated features assessing task success and dialogue costs, and Bodigutla et al. (2019), which proposed domain-independent features that evaluate response quality. However, the growth of LLM-based conversational systems (e.g., ChatGPT, Bing Copilot) means user queries in conversational systems may now be across multiple domains and intents (e.g., task-oriented, QA, chitchat). As such, approaches based on domain-specific features have limited generalizability to these diverse conversational patterns (Deriu et al., 2021).
In this work, we make the key observation that LLMs can achieve both high classification accuracy and fine-grained interpretability at the same time – through their ability to reason about user conversational patterns and identify salient pattern classes that generalize and produce accurate predictions. We propose Supervised Prompting for User satisfaction Rubrics (SPUR). We consider a few-shot scenario, where a small number of training examples are available, and develop a supervised, iterative prompting framework that uses an LLM to (1) extract signals of satisfaction from user utterances in a labeled training set, (2) summarize the reasons into rubrics for identifying satisfaction/dissatisfaction conversational patterns, and (3) apply the rubrics to predict satisfaction labels on unseen conversations.
@@
  • We use knowledge distillation to scale the application of learned rubrics and show the rubrics can continuously improve performance as more training data is available.
@@ 2 Problem Definition and Related Work
Problem Definition. Let a conversation $C$ from session $i$ and consisting of $t$ interaction turns of user-agent utterances be $C_{i}=[U_{1},A_{1},\ldots,U_{t},A_{t}]$. Here $U_{t}$ refers to a user utterance and $A_{t}$ refers to an AI agent utterance. The user-agent utterances $C_{i}$ typically consist of multiple turns, e.g., $t>1$. The conversation also has an overall user satisfaction label $y_{i}\in[-1,+1]$ provided by thumb feedback (e.g., like or dislike).
Our goal is to learn a function $f:C\rightarrow y$ to accurately predict the satisfaction label of unseen conversations and explain the predicted label. In multi-turn conversational sessions, a user can convey their satisfaction (or dissatisfaction) explicitly in their utterances or implicitly through their behavioral interactions with the agent. We refer to these satisfaction/dissatisfaction conversational patterns as SAT/DSAT patterns. Let $\mathcal{S}=\{s_{1},s_{2},\cdots,s_{\infty}\}$ and $\mathcal{D}=\{d_{1},d_{2},\cdots,d_{\infty}\}$ be the set of all interpretable SAT and DSAT patterns respectively. We assume these are latent and unknown. The goal is to identify a subset of SAT/DSAT pattern classes ($\mathcal{S}_{s}\subset\mathcal{S},\mathcal{D}_{s}\subset\mathcal{D}$) that summarize the conversation enough to accurately predict its label: $P(y|C)\approx P\Big{(}y\>\Big{|}\>\mathcal{S}_{s}(C),\mathcal{D}_{s}(C)\Big{)}$.
@@
Related Work. Numerous prior research studies have examined User Satisfaction Evaluation (USE) through the lenses of sentiment analysis (Song et al., 2023, 2019), content analysis (Walker et al., 1997; Sun et al., 2021), and response quality assessment (Schmitt and Ultes, 2015; Bodigutla et al., 2019). While analyzing user sentiment distribution in a dialogue session can enhance the model’s USE capabilities, it is important to note that sentiment analysis is not equivalent to USE (Song et al., 2023). Another common approach involves content analysis, which typically necessitates the employment of human annotators to evaluate interaction quality in a dialogue session (Schmitt and Ultes, 2015; Bodigutla et al., 2019). Afterwards, a classifier is trained to predict user satisfaction based on the features extracted from the annotation process.
@@
Recently, Large Language Models (LLMs) revolutionized the traditional learning framework (Kojima et al., 2022; Wei et al., 2022), especially for natural language processing (NLP) tasks. LLMs have achieved performance comparable to supervised baselines or state-of-the-art results across various NLP tasks with In-Context Learning (ICL). By providing a few examples or hints (Lampinen et al., 2022; Sun et al., 2023) and simple reasoning process (Kojima et al., 2022; Wei et al., 2022), LLMs can provide significant performance boosts for NLP tasks. Hu et al. (2023) further uses LLMs as a user simulator for USE and adopts the user simulator into RLAIF (Bai et al., 2022) for fine-tuning the existing LLM models. For USE with zero-shot prompting (Kojima et al., 2022; Hu et al., 2023), instructions provided by a human may not fit the actual conversation patterns in the data and hence introduce bias. For few-shot prompting Lampinen et al. (2022); Sun et al. (2023), the provided examples are not enough to describe the full distribution of the conversational patterns, and this results in inaccuracies for USE.
@@ 3 SPUR
We propose SPUR for interpretable User Satisfaction Estimation (USE) given a small set of labeled conversation $C$ from a conversational system. Due to the multi-turn and general-purpose nature in such conversational systems, users demonstrate a variety of response patterns when expressing satisfaction or dissatisfaction. Our approach follows the three-phase prompting strategy depicted in Figure 2: Supervised Extraction, Rubric Summarization, and User Satisfaction Estimation. Our three-phase approach is essential for ensuring accuracy, generalization, and interpretability. Through Supervised Extraction, SPUR improves accuracy by capturing the diverse conversational patterns in the training set $\mathcal{C}_{train}=\{C_{1},C_{2},\cdots,C_{N}\}$, which are annotated with thumb feedback. In the Rubric Summarization stage, the LLM improves generalization and interpretability by identifying prominent SAT and DSAT pattern classes among the full set of extracted pattern. Finally, SPUR uses the learned rubrics generated from the previous stage to score user satisfaction on unlabeled conversations. For the ease of understanding, we use mathematical definitions to approximate the process of SPUR in the following three sections.
@@ 3.1 Supervised Extraction
The first step of our framework is Supervised Extraction—where we use a prompt to obtain meaningful and interpretable SAT/DSAT pattern classes from GPT-4, which has an exceptional ability for natural language understanding and reasoning (Ye and Durrett, 2023; Huang et al., 2023; Kojima et al., 2022). Given a conversation $C_{i}$ with its user satisfaction label $y_{i}=+1$, how the user expresses satisfaction in $C_{i}$ can be formulated as: $\widehat{\mathbf{s}}_{i}\approx\underset{s\in\mathcal{S}}{\operatorname{arg\,max_{k}}}\;P(\mathcal{S}|C_{i},y_{i}=+1)$ where $\mathcal{S}=\{s_{1},s_{2},\cdots,s_{\infty}\}$ is the set of all possible SAT patterns. The goal is to identify the top-k potential pattern classes $\widehat{\mathbf{s}}_{i}=\{s_{1},s_{2},\cdots,s_{k}\}\subset\mathcal{S}$ that are exhibited in $C_{i}$ relevant to satisfaction expression. Similarly, $\widehat{\mathbf{d}}_{i}\approx\underset{d\in\mathcal{D}}{\operatorname{arg\,max_{k}}}\;P(\mathcal{D}|C_{i},y_{i}=-1)$, where $\mathcal{D}=\{d_{1},d_{2},\cdots,d_{\infty}\}$ is the set of all possible DSAT patterns.
The prompt for generating the possible $\widehat{\mathbf{s}}$ or $\widehat{\mathbf{d}}$ patterns from $\mathcal{C}_{train}$ is provided in Appendix A.1. In our prompt, we specifically require GPT-4 to restrict $k\leq 3$ for each $C_{i}$. The prompt for DSAT patterns is similar; we only replace “satisfaction” with “dissatisfaction” in the instructions.
For the ease of discussion in the next section, let $\widehat{\mathcal{S}}=\{\widehat{s}_{1},\cdots,\widehat{s}_{N}\}$ denote all the SAT patterns derived from Supervised Extraction and $\widehat{\mathcal{D}}=\{\widehat{d}_{1},\cdots,\widehat{d}_{N}\}$ are all the DSAT patterns.
@@ 3.2 Rubric Summarization
There are two benefits to utilizing the LLM-generated satisfaction and dissatisfaction rubrics from this iterative process. First, the rubrics are developed in a supervised manner from the set of training conversations, $\mathcal{C}_{train}$, thereby ensuring that prominent (and thus predictive) SAT and DSAT pattern classes in the distribution are identified. As a result, the generated rubrics provide a clear guideline for GPT-4 to estimate user satisfaction accurately. Second, the rubrics are generated from more examples than can fit in a single context window. As such, Rubric Summarization improves the generalization for GPT-4 in terms of in-context learning.
@@ 4 Evaluation
4.1 Baselines
We compare SPUR with two LLM-based methods, including ZeroShot and FewShot, and three embedding-based methods: Linear Regression, USDA (Deng et al., 2022) and ASAP (Ye Fanghua, 2023). Note that we choose GPT-4 for all LLM-based methods instead of other smaller language models because smaller language models struggle to accurately generate scores for each rubric item, which results in incorrect SAT scores. The detailed descriptions of the models are as follows:
  1. 1.
    Lin-ada: Linear regression model with ada-002 embedding Ada
  2. 2.
  3. 3.
  4. 4.
    Zero shot: prompt GPT-4 directly to score conversations for user satisfaction with basic reasoning steps by providing explanations.
  5. 5.
    Few shot: prompt GPT-4 directly to score conversations, include 2 examples of labeled conversations to guide GPT-4 to determine user satisfaction and include basic reasoning steps by providing explanations.
  6. 6.
    RQ: prompt GPT-4 with a manually selected features to assess the response quality (Bodigutla et al., 2020) and ask GPT-4 to determine user satisfaction based on the set of features.
  7. 7.
4.2 Dataset
We use four datasets to evaluate the performance of the compared methods. Bing Copilot is a general-purpose and multilingual conversational system, and this dataset includes 50K fully anonymized conversations. [^4] All personal, private, or sensitive information was scrubbed and masked before the conversations were used for this research. The access to the dataset is strictly limited to the authors who conducted hands-on analysis and model development.. MWOZ (Eric et al., 2020), SGD (Rastogi et al., 2020) and ReDial (Siro et al., 2022) are three task-oriented, English conversational systems, and they have 1155, 1638, and 1387 conversations, respectively. These three datasets are further processed and labeled user satisfaction by Sun et al. (2021). Because Sun et al. (2021) labeled user satisfaction by turn, we further process these labels into a label to represent the overall satisfaction of the whole conversation. The preprocessing details are described in Appendix B.
Ethics. As part of the production process, the Bing Copilot data is anonymized, and each conversation is formed by aggregating turns based on a unique conversation ID. Thus, none of the researchers who analyzed the data are able to recover and identify the conversations from any individual user. In addition, this research study was reviewed and approved by representatives from our institutional review board (IRB), as well as our ethics and security teams. No formal IRB certificate was required since we did not conduct any human studies for this paper.
Table 1: Precision (P), Recall (R), and F1 Score (F1) on USE with small training set sizes. The training sizes are shown besides the name of each dataset. The testing size is $80\%$ of the data. The best scores are in bold face.